Each month, a Filipino household generates an average of 28 pounds of newspaper for garbage or recycling. Assume the standard deviation is 2 pounds. If a household is selected at random, find the probability of its generating more than 31 pounds. (

1
Expert's answer
2022-03-21T00:25:24-0400

We have a normal distribution, "\mu=28, \sigma=2."

Let's convert it to the standard normal distribution,

"z=\cfrac{\bar{x}-\mu}{\sigma}" =

"=\cfrac{31-28}{2}=1.5,"

"P(\bar{X}>31)=1-P(\bar{X}<31)=\\
=1-P(Z<1.5)="

"=1-0.9332=0.0668" ​(from z-table)